The length of a rectangle is 8 inches more than its width. The perimeter of the rectangle is 24 inches. Write
and solve an equation to find the width and the length of the rectangle.

Answer by josgarithmetic(39617)

Length, L
width, w

L=w+8 and 2w+2L=24.

2w+2(w+8)=24
w+(w+8)=12
2w+8=12
2w=4
w=2----------------part of answer, width
-
L=w+8=2+8=10 ---------------part of answer, length